# Fast algorithms to improve fair information access in networks

### v1

@@
Abstract
When information spreads across a network via pairwise sharing, large disparities in information access can arise from the network’s structural heterogeneity. Algorithms to improve the fairness of information access seek to maximize the minimum access of a node to information by sequentially selecting new nodes to seed with the spreading information. However, existing algorithms are computationally expensive. Here, we develop and evaluate a set of 10 new scalable algorithms to improve information access in social networks; in order to compare them to the existing state-of-the-art, we introduce both a new performance metric and a new benchmark corpus of networks. Additionally, we investigate the degree to which algorithm performance on minimizing information access gaps can be predicted ahead of time from features of a network’s structure. We find that while no algorithm is strictly superior to all others across networks, our new scalable algorithms are competitive with the state-of-the-art and orders of magnitude faster. We introduce a meta-learner approach that learns which of the fast algorithms is best for a specific network and is on average only 20% less effective than the state-of-the-art performance on held-out data, while about 75-130 times faster. Furthermore, on about 20% of networks the meta-learner’s performance exceeds the state-of-the-art.
I Introduction
Fairness in machine learning is a well studied area, and includes notions of fairness based both on individual characteristics and on group demographics feldman2015certifying ; hardt2016equality ; dwork2012fairness (for surveys, see Refs. caton2024fairness ; pessach2022review ). However, fairness in the context of social networks is comparatively less studied. Individual Fish_2019 and group stoica2019fairness notions of fairness have been introduced that assess individuals’ access to information that spreads across a social network. In this work, we introduce fast algorithms to achieve individual fairness in networks, as defined by Ref. Fish_2019 .
@@ I Introduction
In social networks, access to information is unequally distributed. This inequality is a natural consequence of heterogeneous network connectivity and position—highly-connected or centrally located individuals have more opportunities to receive and spread information cheng2019 ; PICARD201877 ; toth2021 , while peripheral individuals with few connections participate less often in information exchanges PICARD201877 ; Fish_2019 . For example, in a pandemic, access to crucial resources—such as money, food and healthcare—is more difficult for socially disadvantaged groups, in part due to their more limited connectedness in social networks GAROON20081133 . Similarly, connectedness shapes employment opportunities in professional social networks such as LinkedIn: well-connected job seekers are likely to fill lucrative openings sooner than others weaktiesinlinkedin2022 .
The quality of a node’s information access can be formalized mathematically. Given a graph $G=(V,E)$, let $S\subset V$ be the seed set, where $x\in S$ represents an individual who is seeded with information and propagates it across the network via a word-of-mouth process. For a node $i\in V$, we say that the probability $\pi_{i}$ that node $i$ receives the information measures the information access of the individual $i$ under $S$. The notion of individual fairness on networks we will consider Fish_2019 seeks to maximize the minimum such information access $\pi_{i}$ over all individuals $i\in V$, with the aim of focusing interventions on the most disadvantaged with respect to information access in the network.
@@
Figure 1: Algorithm runtime to select 10 new seeds vs. network size for algorithms in Ref. Fish_2019 , averaged over 10 runs on an introduced large set of networks (see Section III.1). Algorithms requiring a Monte Carlo simulation (ProbEst) to select seeds are denoted by a *.
Selecting $k$ nodes at which to seed information to maximize $\min_{i\in V}\pi_{i}$ is NP-hard, and hence past research has proposed heuristics Fish_2019 . However, the heuristics’ performance has been evaluated on a structurally limited range of networks Fish_2019 , and it remains unclear how algorithm performance depends on the structure of the network. Additionally, the algorithms rely on computing $\pi_{i}$ based on the independent cascade model for information propagation, but it is #P-hard to do so exactly provan1983complexity . Therefore, most research has relied on Monte Carlo simulation (here denoted ProbEst; see Section IV.1) to estimate $\pi_{i}$ for each node in the network Fish_2019 . However, ProbEst has high computational cost, making it impractical on very large networks (see Fig. 1). In addition, algorithms that rely on ProbEst can only be evaluated after choosing a value for the parameter $\alpha$, i.e., the independent cascade model’s probability of successful transmission across any given edge kempe_influence_maximization_2003 . The choice of $\alpha$ is not trivial, as the structure of the network influences how well information spreads under a given $\alpha$, and there exists no standard metric for comparing algorithm performance on a given network, or comparing across networks. Finally, it remains unknown whether new, more scalable algorithms can achieve comparable performance to algorithms based on ProbEst, and it is unknown whether one algorithm always generates the best seed sets regardless of the network, and if not, how we might choose the best algorithm given only characteristics of a network.
Progress on these questions would advance our understanding of how to, given a budget constraint and a target network, devise an optimal intervention technique that serves to improve information access of the disadvantaged individuals. This knowledge would be instrumental in designing effective public health campaigns, outreach strategies, etc., that aim to improve fairness in social networks by reaching even the most socially isolated individuals alexander_india_healthcare_2022 . This goal is particularly important because conventional influence maximization techniques have been shown to be ineffective in improving information access of the disadvantaged nodes Fish_2019 .
I.1 Contributions
In our work, we introduce ten new fast algorithms to improve fairness in information access; specifically to maximize the minimum $\pi_{i}$ given a budget of $k$ nodes at which to seed information. We introduce a comprehensive cross-domain network corpus and introduce the concept of spreadability, which we define as the ability of a network to spread information under a given $\alpha$. We subsequently pick $\alpha$ values individually for each network based on a target spreadability. We introduce evaluation mechanisms to assess the quality of fair information access algorithms including a new metric $\beta$, the average marginal increase in access for each additional seed selected. Using the introduced corpus and evaluation mechanisms, we assess the new fast algorithms in comparison to the state-of-the-art. We provide the network corpus and code for this study at https://github.com/rhythmthief/FairnessNetworks. Finally, we introduce a meta-learner approach that chooses the best fast algorithm based on network structure features alone, and is, on average, only 20% less effective than the state-of-the-art while being 75-130 times faster. Furthermore, on about 20% of networks the meta-learner’s performance exceeds the state-of-the-art.
II Related Work
Considerations about the access of individuals to resources in a network build on structural concerns about social networks pioneered by Ref. granovetter1973 . When considering how resources are shared across a network, two popular approaches to spreading information are broadcasting and word-of-mouth NBERw24679 ; kempe_influence_maximization_2003 ; watts_influentials_2007 . Broadcasting is a one-to-many approach, where a single source sends information to all nodes in the network. Word-of-mouth is a many-to-many approach, where information is spread through the network by individuals sharing it with their neighbors NBERw24679 . The latter approach leverages the structure of the network to spread information, and real-world data suggests it can be more effective than broadcasting in certain settings NBERw24679 .
Numerous models of information propagation through networks exist, including the independent cascade model, generalized independent cascade model, and linear threshold model kempe_influence_maximization_2003 ; granovetter1978 ; goldenberg2001 . A popular choice in prior work, also adopted by us, is the independent cascade model, which encapsulates the fundamental dynamics of information spread in networks kempe_influence_maximization_2003 ; Fish_2019 . It operates stochastically, where each edge may facilitate a single attempt to transmit information from an active node to its inactive neighbor, with some probability $\alpha$ of success kempe_influence_maximization_2003 . The independent cascade model’s simplicity provides an effective framework for studying information dissemination across networks.
Under a word-of-mouth approach, information access in a network can be improved via selective seeding, i.e., by choosing an individual or set of individuals to be the first to receive the information NBERw24679 . This idea also underlies the field of influence maximization, which aims to identify the best seed set in order to maximize the spread of information through the network kempe_influence_maximization_2003 . A variety of algorithms have been proposed to solve the influence maximization problem kempe_influence_maximization_2003 ; SINGH20227570 ; Zareie2023 . In contrast to our work, the traditional goal has been to maximize the expected number of activated nodes by selecting a small set of initially activated individuals to seed an information cascade kempe_influence_maximization_2003 . This approach does not account for the fairness of the spread among individuals in the network.
@@ II Related Work
Recently, questions of fairness of information access of individuals and demographic groups in a social network have come to the fore Fish_2019 ; stoica2019fairness ; stoica2020seeding ; ali2021fairness ; mehrotra2022revisiting ; tsang2019group-fairness ; stoica2018algorithmic . Much work focuses on the probability that an individual receives information that spreads through a social network. This work includes studies of seeding information at nodes to improve individual access to that information Fish_2019 , or group access tsang2019group-fairness ; stoica2020seeding ; or interventions to add edges bashardoust2023reducing ; becker2023improving ; bhaskara2024optimizing under varying notions of fairness. Other work considers notions of group fairness based on the network structure jalali2023fairness ; mehrotra2022revisiting ; bashardoust2020information . Ref. saxena2024fairsna provides a recent survey of such work.
@@
Figure 2: Average degree of a network as a function of network size (number of nodes) for the corpus of 174 networks from 6 distinct domains used in our study.
III Network Corpus and Evaluation Methods
III.1 Network Corpus
@@
In order to investigate the effects of network structure on algorithm performance, we construct a corpus of $174$ networks from six domains: biological, social, economic, technological, transportation, and informational. We include non-social networks in our study in order to more fully characterize the behavior of algorithms on structurally diverse real-world networks. When relevant, we report our results by domain, so that social and non-social networks can be contrasted.
The networks in the corpus have at least 500 nodes, and are unipartite, that is, they only have one type of node. We simplified each network by ignoring edge weights, edge direction and self-loops, and took the largest connected component of each network, which itself must contain at least 500 nodes. We also ensured that no domain accounts for more than 25% of all networks in the corpus. Data for the corpus was obtained from  Ghasemian_2019 ; ICON ; Netzschleuder . An overview of the corpus is presented in Figure 2, with summary statistics in Table S1.
III.2 Spreadability
In our work, we leverage the independent cascade model to study the spread of information in networks and estimate $\pi_{i}$. Given a network $G$ and a set of activated nodes $Q\subseteq V$, we grow a forest on $G$ under the independent cascade model by flipping a coin for each edge $e_{ij}$, where $i\in V\backslash Q$, $j\in Q$, exactly once, so that $i$ is added to $Q$ on a successful flip. Each edge $e_{ij}$ is considered as a transmission path at most once, and we define the probability of successful transmission to be $\alpha$. We note that mathematically, it is equivalent to conceptualize the independent cascade as follows: delete every edge in the graph independently with probability $(1-\alpha)$; the set $Q$ of activated nodes is exactly the set of nodes that remain reachable from the seed set $S$. As noted earlier, computing the exact probability of activation for a node $i$, denoted $\pi_{i}$, is #P-hard. As such, we adopt the standard Monte Carlo simulation approach, ProbEst Fish_2019 , but return later to consider practical consequences of this choice.
@@ III.2 Spreadability
Figure 3: Spreadability on a network is quantified by the average fraction of a network’s nodes $\langle|T|\rangle/n$ in a tree $T$ grown through an independent cascade from a random initial seed for a given $\alpha$. We define ‘low’, ‘medium’, and ‘high’ spreadability as the $\alpha$ that activates, on average, $20\%$, $50\%$, and $80\%$ of the network, respectively.

### v2

@@
Abstract.
We consider the problem of selecting $k$ seed nodes in a network to maximize the minimum probability of activation under an independent cascade beginning at these seeds. The motivation is to promote fairness by ensuring that even the least advantaged members of the network have good access to information. Our problem can be viewed as a variant of the classic influence maximization objective, but it appears somewhat more difficult to solve: only heuristics are known. Moreover, the scalability of these methods is sharply constrained by the need to repeatedly estimate access probabilities.
We design and evaluate a suite of $10$ new scalable algorithms which crucially do not require probability estimation. To facilitate comparison with the state-of-the-art, we make three more contributions which may be of broader interest. We introduce a principled method of selecting a pairwise information transmission parameter used in experimental evaluations, as well as a new performance metric which allows for comparison of algorithms across a range of values for the parameter $k$. Finally, we provide a new benchmark corpus of $174$ networks drawn from $6$ domains. Our algorithms retain most of the performance of the state-of-the-art while reducing running time by orders of magnitude. Specifically, a meta-learner approach is on average only $20\%$ less effective than the state-of-the-art on held-out data, but about $75-130$ times faster. Further, the meta-learner’s performance exceeds the state-of-the-art on about $20\%$ of networks, and the magnitude of its running time advantage is maintained on much larger networks.
1. Introduction
Influence maximization (kempe_influence_maximization_2003, ; chen2010scalable, ) is one of the most intensively studied problems in data mining, machine learning, and social network analysis. Given some model of information diffusion, commonly the independent cascade of Kempe, Kleinberg, and Tardos (kempe_influence_maximization_2003, ), the task is to determine where we should seed information such that it spreads as widely as possible. Formally, given a graph $G$ and $S\subseteq V(G)$, for each $i\in V(G)$ let $\pi_{i}(S)$ denote the probability with which $i$ is activated by an independent cascade seeded at $S$; we will just write $\pi_{i}$ when the set $S$ is clear. Then for a given budget $k$, influence maximization is the problem of identifying a set $S\subseteq V(G)$ of cardinality $k$ which maximizes $\sum_{i\in V(G)}\pi_{i}$.
@@ 1. Introduction
Influence maximization is commonly motivated by commercial advertising, e.g., (domingos2001mining, ; richardson2002mining, ; chen2010scalable, ) but is also relevant in various other applications, including public health (leskovec2007cost, ; yadav2016using, ), the distribution of economic opportunities (banerjee2013diffusion, ), and the spread of scientific knowledge (rogers2014diffusion, ). In these applications, the desiderata include not only widespread but also fair dissemination of information. The difficulty of the latter goal is a natural consequence of structural heterogeneity in networks. Highly-connected or centrally located individuals have more opportunities to receive and spread information (cheng2019, ; PICARD201877, ; toth2021, ), while peripheral individuals with few connections participate less often in information exchanges (PICARD201877, ; Fish_2019, ). For example, in a pandemic, access to crucial resources—such as money, food and healthcare—is more difficult for socially disadvantaged groups, in part due to their more limited connectedness in social networks (GAROON20081133, ). Similarly, connectedness shapes employment opportunities in professional social networks such as LinkedIn: well-connected job seekers are likely to fill lucrative openings sooner than others (weaktiesinlinkedin2022, ). This situation has motivated the study of several variations of influence maximization, for example to ensure equitable information access with respect to demographic groups (stoica2018algorithmic, ). In this paper, we focus on the formulation of Fish et al(Fish_2019, ), which adopts a Rawlsian (rawls2009theory, ) notion of fairness in which the goal is to improve the information access of the worst-off individuals. Formally, the problem is as follows:
Maximin Influence Maximization
Input: A graph $G=(V,E)$ and an integer $k$. Task: Select $k$ vertices $S\subseteq V$ maximizing $\min_{i\in V}\pi_{i}$.
This maximin variant is NP-hard, and even a constant-factor approximation would imply P = NP (Fish_2019, ), in stark contrast to the greedy $(1-1/e)$-approximation for the standard objective (kempe_influence_maximization_2003, ). However, Fish et al. show that a heuristic approach (Myopic; see Section 3.1) optimizes the objective well in practice. Indeed, this approach even performs well when evaluated via the classic influence maximization objective, which it does not directly optimize. Unfortunately, this heuristic relies upon repeated Monte Carlo simulations of the activation probabilities $\pi_{i}$, which are a severe computational bottleneck and thereby constrain the scalability of the method; see Figure 1. Additionally, the method has only been evaluated on a small corpus of six networks.
We are thus motivated to (a) introduce new algorithms which alleviate or completely avoid the bottleneck imposed by probability estimation; (b) compare the performance of our methods against that of Myopic in a principled manner,; and (c) expand the number and diversity of networks on which the Maximin Influence Maximization problem has been studied.
@@
Figure 1. Algorithm runtime to select 10 new seeds vs. network size for algorithms in (Fish_2019, ), averaged over 10 runs on an introduced large set of networks (see Section 2.1). Algorithms requiring a Monte Carlo simulation (ProbEst) to select seeds are denoted by a *.
1.1. Related Work and Preliminaries
Fairness in ML and Social Networks.
@@ 1.1. Related Work and Preliminaries
Fairness in machine learning is a well-studied area, and includes notions of fairness based both on individual characteristics and on group demographics (feldman2015certifying, ; hardt2016equality, ; dwork2012fairness, ) (for surveys, see (caton2024fairness, ; pessach2022review, )). Recently, questions of fairness of information access of individuals and demographic groups in a network have come to the fore (Fish_2019, ; stoica2019fairness, ; stoica2020seeding, ; ali2021fairness, ; mehrotra2022revisiting, ; tsang2019group-fairness, ; stoica2018algorithmic, ). A significant body of work focuses on the probability that an individual receives information that spreads through a network. This includes studies of seeding information at nodes to improve individual (Fish_2019, ) or group (tsang2019group-fairness, ; stoica2020seeding, ) access; as well as interventions to add edges (bashardoust2023reducing, ; becker2023improving, ; bhaskara2024optimizing, ) under varying notions of fairness. Other work considers notions of group fairness based on the network structure (jalali2023fairness, ; mehrotra2022revisiting, ; bashardoust2020information, ). Saxena, Fletcher, and Pechenizkiy provide a recent survey of such work (saxena2024fairsna, ). Fair clustering of individuals has also received significant attention, e.g., (chierichetti2017fair, ).
Independent Cascade.
Considerations about the access of individuals to resources in a network build on structural concerns about social networks pioneered by Granovetter (granovetter1973, ). Necessary in any such study is a clear model for the dynamics of information propagation. Numerous models exist, notably including the independent cascade, generalized independent cascade, and linear threshold models (kempe_influence_maximization_2003, ; granovetter1978, ; goldenberg2001, ). A standard choice, also adopted in this paper, is the independent cascade, which can be defined via an iterative spreading process. At each step $i$, some subset $S_{i}$ of nodes is activated, beginning with an initial seed set $S=S_{0}$. In round $i+1$, each edge $uv$ with $u\in S_{i}$, $v\notin\bigcup_{j\leq i}S_{j}$ activates $v$ independently with probability $\alpha$. The process stops after the first round $i$ in which no nodes are activated, i.e., $S_{i}=\emptyset$, and a node $v$ is said to be activated by the cascade if it is activated in any round, i.e., $v\in\bigcup S_{i}$.
Probability Estimation
Important in any analysis of independent cascades is the ability to measure the probability $\pi_{i}$ with which node $i$ is activated. Unfortunately, exact computation of $\pi_{i}$ is #P-hard (chen2010scalable, ). The standard approach is to estimate these values via reverse influence sampling (RIS) (borgs2014maximizing, ; tang2014influence, ), which may also be thought of as performing a series of Monte Carlo simulations of the cascade process. Theoretical bounds on the number of simulations needed to satisfy a given error tolerance $\varepsilon$ have a quadratic dependence on $\varepsilon$ in addition to quasilinear dependence on network size (tang2014influence, ), and so in practice it is common to fix a reasonable number $R$ of simulations, e.g., $R=1000$. Algorithms for influence maximization and related problems then invoke a linear-time (regarding $R$ as a constant) subroutine, referred to here as ProbEst, to provide estimated $\pi_{i}$ values. In practice, however, algorithms requiring such a subroutine can be orders of magnitude slower than those avoiding probability estimation altogether; we again refer to Figure 1. In the influence maximization literature, much effort has been expended to lessen the complexity of ProbEst while retaining quality guarantees, e.g., (tang2015influence, ; nguyen2016stop, ; huang2017revisiting, ), as well as to empirically compare various strategies (ohsaka2020solution, ). In this work, our approach is to develop methods which avoid RIS entirely and evaluate the resulting solution quality via empirical comparison against the state-of-the-art. A purely heuristic approach, while difficult to accept for the classic influence maximization problem, is palatable in our setting because strong inapproximability bounds are known for Maximin Influence Maximization even in the presence of an oracle which perfectly computes the $\pi_{i}$ values (Fish_2019, ).
1.2. Summary of Contributions
Network Corpus.
In Section 2.1, we introduce a large and structurally diverse corpus of $174$ networks on which to benchmark both our algorithms and the state-of-the-art. These networks are drawn from six domains: biological, social, economic, technological, transportation, and informational. Though our primary interest is in social networks, building a corpus from multiple domains enhances the structural diversity of our benchmark set; see the discussion in Section 2.1. Where relevant, we report results by domain.
Spreadability.
When evaluating algorithms in the presence of an independent cascade, it is necessary to choose a value for the parameter $\alpha$ which governs the probability of information transmission along edges. It is common in the literature to choose several values, under the assumption that small values fundamentally represent “low” information spread while large values represent “high” spread. However, this qualitative assessment of a particular $\alpha$ value is not consistent across networks, due to the varying underlying combinatorial structure. We introduce spreadability, which enforces mathematical rigor in the assessment of a particular $\alpha$ value as “low”- or “high”-spreading for a given network. We use this framework to select appropriate $\alpha$ values for each network in our corpus, ensuring that we can measure algorithm performance under multiple distinct regimes of information spread.
Performance Metric.
Even with a fixed $\alpha$ value and a single network, it is non-trivial to compare the performance of algorithms. Among other factors, this is because we must account for varying choices of budget $k$, random number generator seeding, and the impracticality of precisely measuring the problem objective. We introduce a performance metric which incorporates all of these factors, and allows us to evaluate the performance of an algorithm via a single number $\beta$. Intuitively, $\beta$ indicates the marginal improvement in the problem objective that we can expect if we use a given algorithm to add an additional information seed. We defer a more formal description to Section 2.3.
Algorithms.
In Section 3, we introduce $10$ new algorithms for the Maximin Influence Maximization problem, which we partition into three categories. The first two replace the Monte Carlo simulations used by the state-of-the-art to estimate activation probabilities with heuristics based on breadth-first search and personalized page rank. The complexity of our subroutines is similar to the Monte Carlo approach, but the hidden constants are vastly improved, and this is reflected in improved running times. The third category of algorithms eliminates probability estimation entirely, instead using topological features of the network to inform seed selection.
Evaluation, Meta-Learning, and Scalability.
In Section 4, we conduct a comprehensive evaluation of our methods, together with a comparison against the state-of-the-art (Myopic (Fish_2019, )). In Section 4.1, we show that on most networks at least one of our algorithms is nearly as effective as or even more effective than Myopic. Meanwhile, our algorithms are much faster, improving on the running time of Myopic by $1$-$4$ orders of magnitude; see Section 4.2. We build on these insights by introducing a meta-learner (see Section 4.3), which uses structural features of a network to predict which of our algorithms will be most effective on that network. This approach allows us to recover roughly $80\%$ of the performance of Myopic while reducing running time by factors of about $75-130$. Finally, in Section 4.4 we show that these improvements in running time persist even on several much larger networks.
@@ 1.2. Summary of Contributions
Figure 2. Average degree of a network as a function of network size (number of nodes) for the corpus of 174 networks from 6 distinct domains used in our study.
2. Network Corpus and Evaluation Methods
2.1. Network Corpus
@@
In order to investigate the effects of network structure on algorithm performance, we construct a corpus of $174$ networks from six domains: biological (34), social (44), economic (43), technological (32), transportation (17), and informational (4). We include non-social networks in our study in order to more fully characterize the behavior of algorithms on structurally diverse real-world networks. When relevant, we report our results by domain, so that social and non-social networks can be contrasted. An overview of the corpus is presented in Figure 2, with summary statistics in Table S1.
Networks were curated from the Index of Complex Networks (ICON, ), a large-scale index of research-quality networks spanning all domains of science, as well as the Netzschleuder network catalogue and repository (Netzschleuder, ) and the corpus of Ghasemian, Hosseinmardi, and Clauset (Ghasemian_2019, ). All networks included in our corpus are simple graphs, meaning their edges are undirected, unweighted, and there are no self-loops. Further, they are unipartite, i.e., they only have one type of node.
Past work on such corpora indicates that domains (and even subdomains, e.g., online social networks vs. offline social networks) are highly distinguishable based on their structure alone (ikehara2017, ). Hence, a specific effort was made to (i) balance the classes, so that no domain was more than 25% of the corpus, (ii) avoid over-representing networks from particular sources (e.g., Twitter follower networks), and (iii) ensure that the minimum network size was large enough to provide good results for information spreading tasks (minimum number of nodes $n_{\min}=500$). These choices improve the breadth and variety of network structure represented in the corpus, and its utility in the analyses of this study.
@@
Figure 3. Spreadability on a network is quantified by the average fraction of a network’s nodes $\langle|T|\rangle/n$ in a tree $T$ grown through an independent cascade from a random initial seed for a given $\alpha$. We define ‘low’, ‘medium’, and ‘high’ spreadability as the $\alpha$ that activates, on average, $20\%$, $50\%$, and $80\%$ of the network, respectively.
2.2. Spreadability
Background.
In our work, we leverage the independent cascade model (kempe_influence_maximization_2003, ) to study the spread of information in networks and estimate $\pi_{i}$. Given a network $G$ and a set of activated nodes $Q\subseteq V$, we grow a forest on $G$ under the independent cascade model by flipping a coin for each edge $e_{ij}$, where $i\in V\backslash Q$, $j\in Q$, exactly once, so that $i$ is added to $Q$ on a successful flip. Each edge $e_{ij}$ is considered as a transmission path at most once, and we define the probability of successful transmission to be $\alpha$. We note that mathematically, it is equivalent to conceptualize the independent cascade as follows: delete every edge in the graph independently with probability $(1-\alpha)$; the set $Q$ of activated nodes is exactly the set of nodes that remain reachable from the seed set $S$. As noted earlier, computing the exact probability of activation for a node $i$, denoted $\pi_{i}$, is #P-hard (chen2010scalable, ). As such, we adopt the standard Monte Carlo simulation approach, but return later to consider practical consequences of this choice.